solve by facotring (multiple choice question)
9x^3-40x^2-25x=0
choices:
a: 5/9, -5
b. -5/9, 5
c. -5, 0, 5/9
d. -5/9, 0, 5.

Answer by checkley77(12844)

9x^3-40x^2-25x=0
x(9x^2-40x-25)=0
x(9x+5)(x-5)=0
X=0 answer.
9x+5=0
9x=-5
x=-5/9 answer.
x-5=0
x=5 answer.
ANSWER D) IS CORRECT.